OpenStudy (anonymous):

A right prism with a base perimeter of 10 units, a height of 3 units, and a total base area of 7 square units has a surface area of ______ square units.

OpenStudy (anonymous):

S = 2B + Ph (Surface area of a right prism = 2*Area of Base + Perimeter of base * height of prism) S = 2(13) + 6(5) S = 26 + 30 S = 56 units

OpenStudy (anonymous):

thanks. I didn't understand what they were asking first of all, but now i do.

OpenStudy (anonymous):

your welcome :)